question 1. a) How many even numbers can be formed from the digits 1, 2, 3, 4, 5?
b) How many of these numbers are greater than 3000?
Answer #1

Note:Here we are considering that a digit can be used only once; please revert if this is not the condition.

a) as total number of even numbers =N(numbers with 2 in the end)+N(numbers with 4 in the end)=4!+4! =24+24 =48

b)total such numbers =N(first number 3 and last number 2)+N(first number 3 and last number 4)+N(first number 4 and last number 2)+N(first number 5 and last number 2)+N(first number 5 and last number 4)

=3!+3!+3!+3!+3! =6+6+6+6+6=30